OpenStudy (anonymous):

Find a quadratic equation in the form x^2+px+q=0 whose roots are a) 1 and 2 b) 1/2 and 3/4 c) radical 2 -1 nd radical 2 +1 *note : only 2 is under the radical. pleaseeeeeee help

OpenStudy (anonymous):

x^2-(sum of the roots)x+(product of the roots)=0 x^2-(1+2)x+(1*2)=0 x^2-3x+2=0

OpenStudy (anonymous):

Thank you so much